# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 9. Acknowledgments
- Chen et al. (2024b) Yangyi Chen, Karan Sikka, Michael Cogswell, Heng Ji, and Ajay Divakaran. 2024b. Dress: Instructing large vision-language models to align and interact with humans via natural language feedback. In CVPR.
@@
- Chen et al. (2024c) Zhe Chen, Jiannan Wu, Wenhai Wang, Weijie Su, Guo Chen, Sen Xing, Muyan Zhong, Qinglong Zhang, Xizhou Zhu, Lewei Lu, et al. 2024c. Internvl: Scaling up vision foundation models and aligning for generic visual-linguistic tasks. In CVPR.
@@
- Cheng et al. (2019) Shuyu Cheng, Yinpeng Dong, Tianyu Pang, Hang Su, and Jun Zhu. 2019. Improving black-box adversarial attacks with a transfer-based prior. NIPS (2019).
@@
- Cherti et al. (2023) Mehdi Cherti, Romain Beaumont, Ross Wightman, Mitchell Wortsman, Gabriel Ilharco, Cade Gordon, Christoph Schuhmann, Ludwig Schmidt, and Jenia Jitsev. 2023. Reproducible scaling laws for contrastive language-image learning. In CVPR.
@@
- Chiang et al. (2023) Wei-Lin Chiang, Zhuohan Li, Zi Lin, Ying Sheng, Zhanghao Wu, Hao Zhang, Lianmin Zheng, Siyuan Zhuang, Yonghao Zhuang, Joseph E Gonzalez, et al. 2023. Vicuna: An open-source chatbot impressing gpt-4 with 90%* chatgpt quality. (2023). https://lmsys.org/blog/2023-03-30-vicuna/
@@
- Christiano et al. (2017) Paul F Christiano, Jan Leike, Tom Brown, Miljan Martic, Shane Legg, and Dario Amodei. 2017. Deep reinforcement learning from human preferences. NIPS (2017).
@@
- Chua et al. (2009) Tat-Seng Chua, Jinhui Tang, Richang Hong, Haojie Li, Zhiping Luo, and Yantao Zheng. 2009. Nus-wide: a real-world web image database from national university of singapore. In ACM CIVR.
@@
- Cisse et al. (2017) Moustapha Cisse, Yossi Adi, Natalia Neverova, and Joseph Keshet. 2017. Houdini: Fooling deep structured prediction models. arXiv preprint arXiv:1707.05373 (2017).
- Cloud (2023) Alibaba Cloud. 2023. Qwen. (2023). https://tongyi.aliyun.com/qianwen/
@@
- Cordts et al. (2016) Marius Cordts, Mohamed Omran, Sebastian Ramos, Timo Rehfeld, Markus Enzweiler, Rodrigo Benenson, Uwe Franke, Stefan Roth, and Bernt Schiele. 2016. The cityscapes dataset for semantic urban scene understanding. In CVPR.
- Costa et al. (2024) Joana C Costa, Tiago Roxo, Hugo Proença, and Pedro RM Inácio. 2024. How deep learning sees the world: A survey on adversarial attacks & defenses. IEEE Access (2024).
- Croce and Hein (2020a) Francesco Croce and Matthias Hein. 2020a. Minimally distorted adversarial examples with a fast adaptive boundary attack. In ICML.
- Croce and Hein (2020b) Francesco Croce and Matthias Hein. 2020b. Reliable evaluation of adversarial robustness with an ensemble of diverse parameter-free attacks. In ICML.
- Cui et al. (2024) Xuanming Cui, Alejandro Aparcedo, Young Kyun Jang, and Ser-Nam Lim. 2024. On the robustness of large multimodal models against image adversarial attacks. In CVPR.
- Dai et al. (2023) Wenliang Dai, Junnan Li, D Li, AMH Tiong, J Zhao, W Wang, B Li, P Fung, and S Hoi. 2023. Instructblip: Towards general-purpose vision-language models with instruction tuning. arxiv 2023. arXiv preprint arXiv:2305.06500 (2023).
@@
- Deng et al. (2023a) Gelei Deng, Yi Liu, Yuekang Li, Kailong Wang, Ying Zhang, Zefeng Li, Haoyu Wang, Tianwei Zhang, and Yang Liu. 2023a. Jailbreaker: Automated jailbreak across multiple large language model chatbots. arXiv preprint arXiv:2307.08715 (2023).
@@
- Deng et al. (2023b) Yue Deng, Wenxuan Zhang, Sinno Jialin Pan, and Lidong Bing. 2023b. Multilingual jailbreak challenges in large language models. arXiv preprint arXiv:2310.06474 (2023).
@@
- Devlin (2018) Jacob Devlin. 2018. Bert: Pre-training of deep bidirectional transformers for language understanding. arXiv preprint arXiv:1810.04805 (2018).
@@
- Dixon et al. (2018) Lucas Dixon, John Li, Jeffrey Sorensen, Nithum Thain, and Lucy Vasserman. 2018. Measuring and mitigating unintended bias in text classification. In AIES.
@@
- Dong et al. (2023b) Junhao Dong, Yuan Wang, Jianhuang Lai, and Xiaohua Xie. 2023b. Restricted black-box adversarial attack against deepfake face swapping. TIFS (2023).
@@
- Dong et al. (2023a) Yinpeng Dong, Huanran Chen, Jiawei Chen, Zhengwei Fang, Xiao Yang, Yichi Zhang, Yu Tian, Hang Su, and Jun Zhu. 2023a. How Robust is Google’s Bard to Adversarial Image Attacks? arXiv preprint arXiv:2309.11751 (2023).
@@
- Dong et al. (2021) Yinpeng Dong, Shuyu Cheng, Tianyu Pang, Hang Su, and Jun Zhu. 2021. Query-efficient black-box adversarial attacks guided by a transfer-based prior. TPAMI (2021).
@@
- Dong et al. (2018) Yinpeng Dong, Fangzhou Liao, Tianyu Pang, Hang Su, Jun Zhu, Xiaolin Hu, and Jianguo Li. 2018. Boosting adversarial attacks with momentum. In CVPR.
@@
- Dong et al. (2019) Yinpeng Dong, Tianyu Pang, Hang Su, and Jun Zhu. 2019. Evading defenses to transferable adversarial examples by translation-invariant attacks. In CVPR.
@@
- Dong et al. (2017) Yinpeng Dong, Hang Su, Jun Zhu, and Fan Bao. 2017. Towards interpretable deep neural networks by leveraging adversarial examples. arXiv preprint arXiv:1708.05493 (2017).
@@
- Dosovitskiy (2020) Alexey Dosovitskiy. 2020. An image is worth 16x16 words: Transformers for image recognition at scale. arXiv preprint arXiv:2010.11929 (2020).
@@
- Dou et al. (2022) Zi-Yi Dou, Yichong Xu, Zhe Gan, Jianfeng Wang, Shuohang Wang, Lijuan Wang, Chenguang Zhu, Pengchuan Zhang, Lu Yuan, Nanyun Peng, et al. 2022. An empirical study of training end-to-end vision-and-language transformers. In CVPR.
@@
- Du et al. (2021) Zhengxiao Du, Yujie Qian, Xiao Liu, Ming Ding, Jiezhong Qiu, Zhilin Yang, and Jie Tang. 2021. Glm: General language model pretraining with autoregressive blank infilling. arXiv preprint arXiv:2103.10360 (2021).
@@
- Everingham et al. (2010) Mark Everingham, Luc Van Gool, Christopher KI Williams, John Winn, and Andrew Zisserman. 2010. The pascal visual object classes (voc) challenge. IJCV (2010).
@@
- Eykholt et al. (2018) Kevin Eykholt, Ivan Evtimov, Earlence Fernandes, Bo Li, Amir Rahmati, Chaowei Xiao, Atul Prakash, Tadayoshi Kohno, and Dawn Song. 2018. Robust physical-world attacks on deep learning visual classification. In CVPR.
@@
- Face and LAION (2023) Hugging Face and LAION. 2023. Image-aware Decoder Enhanced à la Flamingo with Interleaved Cross-attentionS. (2023). https://huggingface.co/HuggingFaceM4/idefics-9b
@@
- Fang et al. (2023) Yuxin Fang, Wen Wang, Binhui Xie, Quan Sun, Ledell Wu, Xinggang Wang, Tiejun Huang, Xinlong Wang, and Yue Cao. 2023. Eva: Exploring the limits of masked visual representation learning at scale. In CVPR.
- Fawzi et al. (2016) Alhussein Fawzi, Seyed-Mohsen Moosavi-Dezfooli, and Pascal Frossard. 2016. Robustness of classifiers: from adversarial to random noise. NIPS (2016).
@@
- Fawzi et al. (2017) Alhussein Fawzi, Seyed-Mohsen Moosavi-Dezfooli, and Pascal Frossard. 2017. The robustness of deep networks: A geometrical perspective. IEEE Signal Processing Magazine (2017).
@@
- Foret et al. (2020) Pierre Foret, Ariel Kleiner, Hossein Mobahi, and Behnam Neyshabur. 2020. Sharpness-aware minimization for efficiently improving generalization. arXiv preprint arXiv:2010.01412 (2020).
- Fu et al. (2024) Chaoyou Fu, Peixian Chen, Yunhang Shen, Yulei Qin, Mengdan Zhang, Xu Lin, Jinrui Yang, Xiawu Zheng, Ke Li, Xing Sun, Yunsheng Wu, and Rongrong Ji. 2024. MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models. arXiv preprint arXiv:2306.13394 (2024).
@@
- Fu et al. (2023) Xiaohan Fu, Zihan Wang, Shuheng Li, Rajesh K Gupta, Niloofar Mireshghallah, Taylor Berg-Kirkpatrick, and Earlence Fernandes. 2023. Misusing tools in large language models with visual adversarial examples. arXiv preprint arXiv:2310.03185 (2023).
@@
- Ganguli et al. (2022) Deep Ganguli, Liane Lovitt, Jackson Kernion, Amanda Askell, Yuntao Bai, Saurav Kadavath, Ben Mann, Ethan Perez, Nicholas Schiefer, Kamal Ndousse, et al. 2022. Red teaming language models to reduce harms: Methods, scaling behaviors, and lessons learned. arXiv preprint arXiv:2209.07858 (2022).
@@
- Gao et al. (2024a) Kuofeng Gao, Yang Bai, Jiawang Bai, Yong Yang, and Shu-Tao Xia. 2024a. Adversarial robustness for visual grounding of multimodal large language models. arXiv preprint arXiv:2405.09981 (2024).
@@
- Gao et al. (2024b) Kuofeng Gao, Yang Bai, Jindong Gu, Shu-Tao Xia, Philip Torr, Zhifeng Li, and Wei Liu. 2024b. Inducing high energy-latency of large vision-language models with verbose images. arXiv preprint arXiv:2401.11170 (2024).
@@
- Gao et al. (2020a) Lianli Gao, Qilong Zhang, Jingkuan Song, and Heng Tao Shen. 2020a. Patch-wise++ perturbation for adversarial targeted attacks. arXiv preprint arXiv:2012.15503 (2020).
- Gao et al. (2023) Peng Gao, Jiaming Han, Renrui Zhang, Ziyi Lin, Shijie Geng, Aojun Zhou, Wei Zhang, Pan Lu, Conghui He, Xiangyu Yue, et al. 2023. Llama-adapter v2: Parameter-efficient visual instruction model. arXiv preprint arXiv:2304.15010 (2023).
- GeekPwn (2018) GeekPwn. 2018. GeekPwn CAAD 2018. (2018). https://en.caad.geekpwn.org/
- Gehman et al. (2020) Samuel Gehman, Suchin Gururangan, Maarten Sap, Yejin Choi, and Noah A Smith. 2020. Realtoxicityprompts: Evaluating neural toxic degeneration in language models. arXiv preprint arXiv:2009.11462 (2020).
- Girdhar et al. (2023) Rohit Girdhar, Alaaeldin El-Nouby, Zhuang Liu, Mannat Singh, Kalyan Vasudev Alwala, Armand Joulin, and Ishan Misra. 2023. Imagebind: One embedding space to bind them all. In CVPR.
- Girshick et al. (2014) Ross Girshick, Jeff Donahue, Trevor Darrell, and Jitendra Malik. 2014. Rich feature hierarchies for accurate object detection and semantic segmentation. In CVPR.
- Gondim-Ribeiro et al. (2018) George Gondim-Ribeiro, Pedro Tabacof, and Eduardo Valle. 2018. Adversarial attacks on variational autoencoders. arXiv preprint arXiv:1806.04646 (2018).
- Gong et al. (2023a) Tao Gong, Chengqi Lyu, Shilong Zhang, Yudong Wang, Miao Zheng, Qian Zhao, Kuikun Liu, Wenwei Zhang, Ping Luo, and Kai Chen. 2023a. Multimodal-gpt: A vision and language model for dialogue with humans. arXiv preprint arXiv:2305.04790 (2023).
- Gong et al. (2023b) Yichen Gong, Delong Ran, Jinyuan Liu, Conglei Wang, Tianshuo Cong, Anyu Wang, Sisi Duan, and Xiaoyun Wang. 2023b. Figstep: Jailbreaking large vision-language models via typographic visual prompts. arXiv preprint arXiv:2311.05608 (2023).
- Goodfellow et al. (2020) Ian Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron Courville, and Yoshua Bengio. 2020. Generative adversarial networks. Commun. ACM (2020).
- Goodfellow et al. (2014) Ian J Goodfellow, Jonathon Shlens, and Christian Szegedy. 2014. Explaining and harnessing adversarial examples. arXiv preprint arXiv:1412.6572 (2014).
- Google (2023) Google. 2023. Bard (Gemini). (2023). https://gemini.google.com/
- Google (2024) Google. 2024. Gemini usage policies. (2024). https://ai.google.dev/gemini-api/docs/safety-settings
- Gou et al. (2024) Yunhao Gou, Kai Chen, Zhili Liu, Lanqing Hong, Hang Xu, Zhenguo Li, Dit-Yan Yeung, James T Kwok, and Yu Zhang. 2024. Eyes closed, safety on: Protecting multimodal llms via image-to-text transformation. arXiv preprint arXiv:2403.09572 (2024).
- Gowal et al. (2019) Sven Gowal, Jonathan Uesato, Chongli Qin, Po-Sen Huang, Timothy Mann, and Pushmeet Kohli. 2019. An alternative surrogate loss for pgd-based adversarial testing. arXiv preprint arXiv:1910.09338 (2019).
- Goyal et al. (2017) Yash Goyal, Tejas Khot, Douglas Summers-Stay, Dhruv Batra, and Devi Parikh. 2017. Making the v in vqa matter: Elevating the role of image understanding in visual question answering. In CVPR.
- gpt 4o (2023) Hello gpt 4o. 2023. OpenAI. (2023). https://openai.com/index/hello-gpt-4o/
@@
- Gu et al. (2024) Xiangming Gu, Xiaosen Zheng, Tianyu Pang, Chao Du, Qian Liu, Ye Wang, Jing Jiang, and Min Lin. 2024. Agent smith: A single image can jailbreak one million multimodal llm agents exponentially fast. arXiv preprint arXiv:2402.08567 (2024).
- Guo et al. (2018) Chuan Guo, Jared S Frank, and Kilian Q Weinberger. 2018. Low frequency adversarial perturbation. arXiv preprint arXiv:1809.08758 (2018).
- Guo et al. (2017) Chuan Guo, Mayank Rana, Moustapha Cisse, and Laurens Van Der Maaten. 2017. Countering adversarial images using input transformations. arXiv preprint arXiv:1711.00117 (2017).
- Guo et al. (2023) Jiaxian Guo, Junnan Li, Dongxu Li, Anthony Meng Huat Tiong, Boyang Li, Dacheng Tao, and Steven Hoi. 2023. From images to textual prompts: Zero-shot visual question answering with frozen large language models. In CVPR.
- Guo et al. (2024) Qi Guo, Shanmin Pang, Xiaojun Jia, and Qing Guo. 2024. Efficiently Adversarial Examples Generation for Visual-Language Models under Targeted Transfer Scenarios using Diffusion Models. arXiv preprint arXiv:2404.10335 (2024).
- Hao et al. (2019) Yaru Hao, Li Dong, Furu Wei, and Ke Xu. 2019. Visualizing and understanding the effectiveness of BERT. arXiv preprint arXiv:1908.05620 (2019).
- Hasan et al. (2024) Adib Hasan, Ileana Rugina, and Alex Wang. 2024. Pruning for protection: Increasing jailbreak resistance in aligned llms without fine-tuning. arXiv preprint arXiv:2401.10862 (2024).
@@
- Hessel et al. (2021) Jack Hessel, Ari Holtzman, Maxwell Forbes, Ronan Le Bras, and Yejin Choi. 2021. Clipscore: A reference-free evaluation metric for image captioning. arXiv preprint arXiv:2104.08718 (2021).
- Heusel et al. (2017) Martin Heusel, Hubert Ramsauer, Thomas Unterthiner, Bernhard Nessler, and Sepp Hochreiter. 2017. Gans trained by a two time-scale update rule converge to a local nash equilibrium. NIPS (2017).
- Hoffmann et al. (2022) Jordan Hoffmann, Sebastian Borgeaud, Arthur Mensch, Elena Buchatskaya, Trevor Cai, Eliza Rutherford, Diego de Las Casas, Lisa Anne Hendricks, Johannes Welbl, Aidan Clark, et al. 2022. Training compute-optimal large language models. arXiv preprint arXiv:2203.15556 (2022).
@@
- Huang et al. (2008) Gary B Huang, Marwan Mattar, Tamara Berg, and Eric Learned-Miller. 2008. Labeled faces in the wild: A database forstudying face recognition in unconstrained environments. In Workshop on faces in’Real-Life’Images: detection, alignment, and recognition.
- Huang et al. (2023a) Hao Huang, Ziyan Chen, Huanran Chen, Yongtao Wang, and Kevin Zhang. 2023a. T-sea: Transfer-based self-ensemble attack on object detection. In CVPR.
- Huang et al. (2019) Qian Huang, Isay Katsman, Horace He, Zeqi Gu, Serge Belongie, and Ser-Nam Lim. 2019. Enhancing adversarial example transferability with an intermediate level attack. In ICCV.
- Huang et al. (2017) Sandy Huang, Nicolas Papernot, Ian Goodfellow, Yan Duan, and Pieter Abbeel. 2017. Adversarial attacks on neural network policies. arXiv preprint arXiv:1702.02284 (2017).
- Huang et al. (2024) Yao Huang, Yinpeng Dong, Shouwei Ruan, Xiao Yang, Hang Su, and Xingxing Wei. 2024. Towards Transferable Targeted 3D Adversarial Attack in the Physical World. In CVPR.
- Huang et al. (2023b) Yangsibo Huang, Samyak Gupta, Mengzhou Xia, Kai Li, and Danqi Chen. 2023b. Catastrophic jailbreak of open-source llms via exploiting generation. arXiv preprint arXiv:2310.06987 (2023).

### v2

@@ 9. Acknowledgments
- Chen et al. (2024a) Y Chen et al. 2024a. Dress: Instructing large vision-language models to align and interact with humans via natural language feedback. In CVPR.
- Chen et al. (2023c) Yang Chen, Ethan Mendes, et al. 2023c. Can language models be instructed to protect personal information? arXiv:2310.02224.
@@
- Chen et al. (2024c) Zhe Chen, Jiannan Wu, et al. 2024c. Internvl: Scaling up vision foundation models and aligning for generic visual-linguistic tasks. In CVPR.
@@
- Cheng et al. (2019) Shuyu Cheng, Yinpeng Dong, Tianyu Pang, Hang Su, et al. 2019. Improving black-box adversarial attacks with a transfer-based prior. In NIPS.
@@
- Cherti et al. (2023) Mehdi Cherti, Romain Beaumont, Ross Wightman, et al. 2023. Reproducible scaling laws for contrastive language-image learning. In CVPR.
@@
- Chiang et al. (2023) Wei-Lin Chiang, Zhuohan Li, Zi Lin, Ying Sheng, Zhanghao Wu, Hao Zhang, Lianmin Zheng, Siyuan Zhuang, Yonghao Zhuang, Joseph E Gonzalez, et al. 2023. Vicuna: An open-source chatbot impressing gpt-4 with 90%* chatgpt quality. https://lmsys.org/blog/2023-03-30-vicuna/
@@
- Christiano et al. (2017) Paul F Christiano, Jan Leike, Tom Brown, Miljan Martic, et al. 2017. Deep reinforcement learning from human preferences. In NIPS.
@@
- Chua et al. (2009) Tat-Seng Chua, Jinhui Tang, et al. 2009. Nus-wide: a real-world web image database from national university of singapore. In ACM CIVR.
@@
- Cisse et al. (2017) Moustapha Cisse, Yossi Adi, Natalia Neverova, et al. 2017. Houdini: Fooling deep structured prediction models. arXiv:1707.05373.
- Clements and Lao (2022) Joseph Clements and Yingjie Lao. 2022. In pursuit of preserving the fidelity of adversarial images. In ICASSP.
- Cloud (2023) Alibaba Cloud. 2023. Qwen. https://tongyi.aliyun.com/qianwen/
@@
- Cordts et al. (2016) Marius Cordts, Mohamed Omran, Sebastian Ramos, et al. 2016. The cityscapes dataset for semantic urban scene understanding. In CVPR.
- Costa et al. (2024) Joana C Costa, Tiago Roxo, et al. 2024. How deep learning sees the world: A survey on adversarial attacks & defenses. IEEE Access.
- Croce et al. (2020) Francesco Croce et al. 2020. Reliable evaluation of adversarial robustness with an ensemble of diverse parameter-free attacks. In ICML.
- Croce and Hein (2020) Francesco Croce and Matthias Hein. 2020. Minimally distorted adversarial examples with a fast adaptive boundary attack. In ICML.
- Cui et al. (2024) Xuanming Cui, Alejandro Aparcedo, et al. 2024. On the robustness of large multimodal models against image adversarial attacks. In CVPR.
- Dai et al. (2023) Wenliang Dai, Junnan Li, et al. 2023. Instructblip: Towards general-purpose vision-language models with instruction tuning. arXiv:2305.06500.
@@
- Deng et al. (2023a) Gelei Deng, Yi Liu, et al. 2023a. Jailbreaker: Automated jailbreak across multiple large language model chatbots. arXiv:2307.08715.
@@
- Deng et al. (2023b) Yue Deng, Wenxuan Zhang, Sinno Jialin Pan, et al. 2023b. Multilingual jailbreak challenges in large language models. arXiv:2310.06474.
@@
- Devlin (2018) Jacob Devlin. 2018. Bert: Pre-training of deep bidirectional transformers for language understanding. arXiv:1810.04805.
@@
- Dixon et al. (2018) Lucas Dixon, John Li, Jeffrey Sorensen, Nithum Thain, et al. 2018. Measuring and mitigating unintended bias in text classification. In AIES.
@@
- Dong et al. (2023b) Junhao Dong, Yuan Wang, Jianhuang Lai, et al. 2023b. Restricted black-box adversarial attack against deepfake face swapping. TIFS.
@@
- Dong et al. (2023a) Yinpeng Dong, Huanran Chen, Jiawei Chen, et al. 2023a. How Robust is Google’s Bard to Adversarial Image Attacks? arXiv:2309.11751.
@@
- Dong et al. (2021) Yinpeng Dong, Shuyu Cheng, et al. 2021. Query-efficient black-box adversarial attacks guided by a transfer-based prior. TPAMI.
@@
- Dong et al. (2018) Yinpeng Dong, Fangzhou Liao, Tianyu Pang, Hang Su, Jun Zhu, Xiaolin Hu, et al. 2018. Boosting adversarial attacks with momentum. In CVPR.
@@
- Dong et al. (2019) Yinpeng Dong, Tianyu Pang, Hang Su, et al. 2019. Evading defenses to transferable adversarial examples by translation-invariant attacks. In CVPR.
@@
- Dong et al. (2017) Yinpeng Dong, Hang Su, et al. 2017. Towards interpretable deep neural networks by leveraging adversarial examples. arXiv:1708.05493.
@@
- Dosovitskiy (2020) Alexey Dosovitskiy. 2020. An image is worth 16x16 words: Transformers for image recognition at scale. arXiv:2010.11929.
@@
- Dou et al. (2022) Zi-Yi Dou, Yichong Xu, Zhe Gan, et al. 2022. An empirical study of training end-to-end vision-and-language transformers. In CVPR.
@@
- Du et al. (2021) Zhengxiao Du, Yujie Qian, et al. 2021. Glm: General language model pretraining with autoregressive blank infilling. arXiv:2103.10360.
@@
- Everingham et al. (2010) Mark Everingham, Luc Van Gool, Christopher KI Williams, et al. 2010. The pascal visual object classes (voc) challenge. IJCV.
@@
- Eykholt et al. (2018) Kevin Eykholt, Ivan Evtimov, Earlence Fernandes, Bo Li, et al. 2018. Robust physical-world attacks on deep learning visual classification. In CVPR.
@@
- Face and LAION (2023) Hugging Face and LAION. 2023. Image-aware Decoder Enhanced à la Flamingo with Interleaved Cross-attentionS. https://huggingface.co/HuggingFaceM4/idefics-9b
@@
- Fang et al. (2023) Yuxin Fang, Wen Wang, Binhui Xie, Quan Sun, et al. 2023. Eva: Exploring the limits of masked visual representation learning at scale. In CVPR.
@@
- Fawzi et al. (2017) Alhussein Fawzi et al. 2017. The robustness of deep networks: A geometrical perspective. IEEE Signal Processing Magazine.
- Fawzi et al. (2016) Alhussein Fawzi, Seyed-Mohsen Moosavi-Dezfooli, et al. 2016. Robustness of classifiers: from adversarial to random noise. In NIPS.
@@
- Foret et al. (2020) Pierre Foret, Ariel Kleiner, et al. 2020. Sharpness-aware minimization for efficiently improving generalization. arXiv:2010.01412.
- Fredrikson et al. (2015) Matt Fredrikson et al. 2015. Model inversion attacks that exploit confidence information and basic countermeasures. In ACM SIGSAC.
- Fu et al. (2024) Chaoyou Fu et al. 2024. MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models. arXiv:2306.13394.
@@
- Fu et al. (2023) Xiaohan Fu, Zihan Wang, et al. 2023. Misusing tools in large language models with visual adversarial examples. arXiv:2310.03185.
@@
- Ganguli et al. (2022) Deep Ganguli et al. 2022. Red teaming language models to reduce harms: Methods, scaling behaviors, and lessons learned. arXiv:2209.07858.
@@
- Gao et al. (2024a) Kuofeng Gao, Yang Bai, et al. 2024a. Adversarial robustness for visual grounding of multimodal large language models. arXiv:2405.09981.
@@
- Gao et al. (2024b) Kuofeng Gao, Yang Bai, et al. 2024b. Inducing high energy-latency of large vision-language models with verbose images. arXiv:2401.11170.
- Gao et al. (2020a) Lianli Gao, Qilong Zhang, Jingkuan Song, et al. 2020a. Patch-wise++ perturbation for adversarial targeted attacks. arXiv:2012.15503.
@@
- Gao et al. (2023a) Peng Gao, Jiaming Han, et al. 2023a. Llama-adapter v2: Parameter-efficient visual instruction model. arXiv:2304.15010.
- Gao et al. (2023b) Yinghua Gao, Dongxian Wu, Jingfeng Zhang, et al. 2023b. On the effectiveness of adversarial training against backdoor attacks. TNNLS.
- GeekPwn (2018) GeekPwn. 2018. GeekPwn CAAD 2018. https://en.caad.geekpwn.org/
- Gehman et al. (2020) Samuel Gehman et al. 2020. Realtoxicityprompts: Evaluating neural toxic degeneration in language models. arXiv:2009.11462.
- Girdhar et al. (2023) Rohit Girdhar, Alaaeldin El-Nouby, Zhuang Liu, Mannat Singh, et al. 2023. Imagebind: One embedding space to bind them all. In CVPR.
- Girshick et al. (2014) Ross Girshick, Jeff Donahue, et al. 2014. Rich feature hierarchies for accurate object detection and semantic segmentation. In CVPR.
- Goldwasser et al. (2020) Shafi Goldwasser, Adam Tauman Kalai, et al. 2020. Beyond perturbations: Learning guarantees with arbitrary adversarial test examples. In NIPS.
- Gondim-Ribeiro et al. (2018) George Gondim-Ribeiro, Pedro Tabacof, and Eduardo Valle. 2018. Adversarial attacks on variational autoencoders. arXiv:1806.04646.
- Gong et al. (2023a) Tao Gong, Chengqi Lyu, et al. 2023a. Multimodal-gpt: A vision and language model for dialogue with humans. arXiv:2305.04790.
- Gong et al. (2023b) Yichen Gong, Delong Ran, et al. 2023b. Figstep: Jailbreaking large vision-language models via typographic visual prompts. arXiv:2311.05608.
- Goodfellow et al. (2020) Ian Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, et al. 2020. Generative adversarial networks. Commun. ACM.
- Goodfellow et al. (2014) Ian J Goodfellow, Jonathon Shlens, and Christian Szegedy. 2014. Explaining and harnessing adversarial examples. arXiv:1412.6572.
- Google (2023) Google. 2023. Bard (Gemini). https://gemini.google.com/
- Google (2024) Google. 2024. Gemini usage policies. https://ai.google.dev/gemini-api/docs/safety-settings
- Gou et al. (2024) Yunhao Gou, Kai Chen, et al. 2024. Eyes closed, safety on: Protecting multimodal llms via image-to-text transformation. arXiv:2403.09572.
- Gowal et al. (2019) Sven Gowal, Jonathan Uesato, et al. 2019. An alternative surrogate loss for pgd-based adversarial testing. arXiv:1910.09338.
- Goyal et al. (2017) Yash Goyal, Tejas Khot, et al. 2017. Making the v in vqa matter: Elevating the role of image understanding in visual question answering. In CVPR.
- gpt 4o (2023) Hello gpt 4o. 2023. OpenAI. https://openai.com/index/hello-gpt-4o/
@@
- Gu et al. (2022) Jindong Gu et al. 2022. Segpgd: An effective and efficient adversarial attack for evaluating and boosting segmentation robustness. In ECCV.
- Gu et al. (2024) Xiangming Gu et al. 2024. Agent smith: A single image can jailbreak one million multimodal llm agents exponentially fast. arXiv:2402.08567.
- Guo et al. (2018) Chuan Guo, Jared S Frank, and Kilian Q Weinberger. 2018. Low frequency adversarial perturbation. arXiv:1809.08758.
- Guo et al. (2017) Chuan Guo, Mayank Rana, et al. 2017. Countering adversarial images using input transformations. arXiv:1711.00117.
- Guo et al. (2023) Jiaxian Guo et al. 2023. From images to textual prompts: Zero-shot visual question answering with frozen large language models. In CVPR.
- Guo et al. (2024) Qi Guo, Shanmin Pang, Xiaojun Jia, and Qing Guo. 2024. Efficiently Adversarial Examples Generation for Visual-Language Models under Targeted Transfer Scenarios using Diffusion Models. arXiv:2404.10335.
- Hameed and Gyorgy (2021) Muhammad Zaid Hameed and Andras Gyorgy. 2021. Perceptually constrained adversarial attacks. arXiv:2102.07140.
- Hao et al. (2019) Yaru Hao, Li Dong, Furu Wei, and Ke Xu. 2019. Visualizing and understanding the effectiveness of BERT. arXiv:1908.05620.
- Hasan et al. (2024) Adib Hasan et al. 2024. Pruning for protection: Increasing jailbreak resistance in aligned llms without fine-tuning. arXiv:2401.10862.
@@
- Hessel et al. (2021) Jack Hessel, Ari Holtzman, et al. 2021. Clipscore: A reference-free evaluation metric for image captioning. arXiv:2104.08718.
- Heusel et al. (2017) Martin Heusel, Hubert Ramsauer, et al. 2017. Gans trained by a two time-scale update rule converge to a local nash equilibrium. In NIPS.
- Hoffmann et al. (2022) Jordan Hoffmann, Sebastian Borgeaud, et al. 2022. Training compute-optimal large language models. arXiv:2203.15556.
- Hönig et al. (2024) Robert Hönig, Javier Rando, et al. 2024. Adversarial perturbations cannot reliably protect artists from generative ai. arXiv:2406.12027.
@@
- Huang et al. (2008) Gary B Huang et al. 2008. Labeled faces in the wild: A database forstudying face recognition in unconstrained environments. In ECCV Workshop.
- Huang et al. (2023a) Hao Huang, Ziyan Chen, Huanran Chen, et al. 2023a. T-sea: Transfer-based self-ensemble attack on object detection. In CVPR.
- Huang et al. (2019) Qian Huang, Isay Katsman, Horace He, et al. 2019. Enhancing adversarial example transferability with an intermediate level attack. In ICCV.
- Huang et al. (2017) Sandy Huang, Nicolas Papernot, Ian Goodfellow, et al. 2017. Adversarial attacks on neural network policies. arXiv:1702.02284.
- Huang et al. (2020) W Ronny Huang, Jonas Geiping, Liam Fowl, et al. 2020. Metapoison: Practical general-purpose clean-label data poisoning. In NIPS.
- Huang et al. (2023b) Yangsibo Huang et al. 2023b. Catastrophic jailbreak of open-source llms via exploiting generation. arXiv:2310.06987.
- Huang et al. (2024) Yao Huang, Yinpeng Dong, Shouwei Ruan, et al. 2024. Towards Transferable Targeted 3D Adversarial Attack in the Physical World. In CVPR.